kgfok26:

If the total operating cost for a 20-acre organic strawberry farm is $551,520, how many acres are on a farm with a total operating cost of $441,216?

kittybasil:

How much is one acre?\[\frac{$551,520}{20\text{ acres}}\]This is your conversion factor. Please note that it can be flipped over if you need to, in order to cancel out units.


You have a farm that costs $441,216. We want to know how many acres it is, so flip your conversion factor and get the cost units on the bottom. \[\large{$441,216\times\frac{20\text{ acres}}{$551,520}}\]\[\large{=\color{red}{\frac{441,216\times20}{551,520}}\text{ acres}}\]
Simplify the fraction from the last part (in RED) and you should have your answer.